An alternating pd 220V amplitude and frequncy 50Hz is applied to a series circuit containing R=400 ohms, L=0.35H and C= 18 micro farad (a) what is Z? (b) what is the current amplitude ,is it lagging the supply pd (c) what is the power factor?



1
Expert's answer
2021-02-12T06:10:53-0500

a)

The impedance of the circuit is

"Z=\sqrt{R^2+(2\pi fL-\frac{1}{2 \pi f C})^2}"

"=\sqrt{(400)^2+(2 \pi (50)(0.35)-\frac{1}{2 \pi (50) (18 \times 10^{-6})})^2}"

="405.5 \Omega"

b)

The current amplitude is

"I=\frac{V}{Z}=\frac{220}{405.5}=0.54 A"

c)

The power factor is

"=\frac{R}{Z}=\frac{400}{405.5}=0.986"